OpenStudy (anonymous):

At a school football game, an adult ticket costs $4 and a student ticket costs $1. A total of 250 people attend the game, earning the school $427. How many students attended the game?

OpenStudy (anonymous):

Let A = the number of adults; Let S = the number of students So $4A + $1S = $427 A + S = 250 So A= 250 - S Substitute into the first equation $4(250-S) + $1S = $427 So S = 191 students